OpenStudy (anonymous):

•Use the two points on the line and the point-slope formula, y - y1 = m(x - x1), to write the equation of the line.

OpenStudy (anonymous):

58,69/61,70

OpenStudy (anonymous):

Do you know how to find the slope of the line containing those two points?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok, what do you get for that?

OpenStudy (anonymous):

y - y1 = m(x - x1 58,69/61,70

OpenStudy (anonymous):

The slope of a line containing two points is m = (y2 - y1) / (x2 - x1) m = (70 - 69) / ( 61 - 58) m = 1 / 3

OpenStudy (anonymous):

thank u ser

OpenStudy (anonymous):

np, know how to finish it?

OpenStudy (anonymous):

what can u help with one more

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

4.Using the equation that you found in question 2, approximately how tall is a person whose arm span is 66 inches?

OpenStudy (anonymous):

Plug the value that @Malaria got for the slope into the point-slope equation for m. Then pick ONE of the two points you provided. Just ONE. And plug in the respect x and y values into the point-slope equation for x1 and y1.

OpenStudy (anonymous):

o ok

OpenStudy (anonymous):

respect = respective. keyboard's stuck on slow today.

OpenStudy (anonymous):

which point did you pick?

OpenStudy (anonymous):

58,69/61,70

OpenStudy (anonymous):

which? (58, 69) or (61, 70)? You get to decide which. There is no "wrong" answer in this step of the process.

OpenStudy (anonymous):

58,69

OpenStudy (anonymous):

Cool. So (58, 69) = (x1, y1). And 1/3 = m. We'll plug those values into the point-slope form equation. y - 69 = 1/3(x - 58)

OpenStudy (anonymous):

i get it now

OpenStudy (anonymous):

Cool.\[y = (1/3)x + (149/3)\]